OpenStudy (anonymous):

Given the functions f(x) = √x-8 and g(x) = 4/2x+1 a) Find (fog)(x) b) Find g^-1(f^-1(x))

OpenStudy (anonymous):

@ybarrap

OpenStudy (ybarrap):

try it and I'll check

OpenStudy (ybarrap):

be back in a few minutes after you try...

OpenStudy (anonymous):

\[fog(x)=\frac{ 4 }{2(\sqrt{x+8)+1} ? }\]

OpenStudy (anonymous):

second part I still don't understand if you could explain to me please how the inverse function will work

OpenStudy (anonymous):

@freckles

OpenStudy (ybarrap):

It might be easier if you write $$ (f\circ g)(x) \text{ as } f(g(x)) $$ so $$ f(g(x))=f\left (\cfrac{4}{2x}+1\right )=\sqrt{g(x)-8}=\sqrt{\left (\cfrac{ 4}{2x}+1\right )-8} $$ I'll let you simplify. Next, for the inverse compositions Check that the following are the inverses of f and g: $$ f^{-1}(x)=x^2+8\\ g^{-1}(x)=\cfrac{2}{x-1}\\ $$ Then $$ \left (g^{-1}\circ f^{-1}\right )(x)=g^{-1}\left (f^{-1}(x)\right )\\ =g^{-1}\left (x^2+8\right )=\cfrac{2}{f^{-1}(x)-1}=\cfrac{2}{(x^2+8)-1} $$ Simplify and you're done! Make sense?